Question 9 4 pts Solve the recurrence relation an an-1+2 with a = 4 (Hint: This will telescope.)
Answer #1

an= an-, tan with ao=4 an- anni = 2 anria an-2= ahot an-2-an-z = 2n-2 add » az -a, .= 2² a, -a. = an-ao = 2h + g h -1 + g h